Rectangular stage. One side of a rectangular stage is 2 meters longer than the other. If the diagnonal is 10 meters, then what are the lengths of the sides?
Answer by checkley71(8403)

x^2+(x+2)^2=10^2
x^2+x^2+4x+4=100
2x^2+4x+4-100=0
2x^2+4x-96=0
2(x^2+2x-48)=0
2(x+8)(x-6)=0
x-6=0
x=6 for one side
6+2=8 for the other side.
proof.
6^2+8^2=100
36+64=100
100=100